OpenStudy (anonymous):

Solve the system by substitution. {2x+y= -11} {3x-4y=11}

OpenStudy (anonymous):

Multiply the first equation by 4

OpenStudy (anonymous):

We get 8x+4y = -44

OpenStudy (anonymous):

Add both equations

OpenStudy (anonymous):

We get 11x=-33 x=-3

OpenStudy (anonymous):

Substitute c= -3 in the first equation 2(-3)+y=-11 -6+y=-11 y=-11+6 y=-5

OpenStudy (anonymous):

U got it?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

Medal?